OpenStudy (anonymous):

Solve for x: 1 over 2 x + 2 = 22.

OpenStudy (anonymous):

subtract 2 from both sides: 2x+2-2 = 22-2 2x=20 x =??

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

\[\frac{1}{2x}+2=22\]\[\frac{1}{2x+2}=22\]Which problem is yours?